### v1

@@ Abstract
Achieving a balance between lightweight design and high performance remains a significant challenge for speech enhancement (SE) tasks on resource-constrained devices. Existing state-of-the-art methods, such as MUSE, have established a strong baseline with only 0.51M parameters by introducing a Multi-path Enhanced Taylor (MET) transformer and Deformable Embedding (DE). However, an in-depth analysis reveals that MUSE still suffers from efficiency bottlenecks: the MET module relies on a complex "approximate-compensate" mechanism to mitigate the limitations of Taylor-expansion-based attention, while the offset calculation for deformable embedding introduces additional computational burden. This paper proposes IMSE, a systematically optimized and ultra-lightweight network. We introduce two core innovations: 1) Replacing the MET module with Amplitude-Aware Linear Attention (MALA). MALA fundamentally rectifies the "amplitude-ignoring" problem in linear attention by explicitly preserving the norm information of query vectors in the attention calculation, achieving efficient global modeling without an auxiliary compensation branch. 2) Replacing the DE module with Inception Depthwise Convolution (IDConv). IDConv borrows the Inception concept, decomposing large-kernel operations into efficient parallel branches (square, horizontal, and vertical strips), thereby capturing spectrogram features with extremely low parameter redundancy. Extensive experiments on the VoiceBank+DEMAND dataset demonstrate that, compared to the MUSE baseline, IMSE significantly reduces the parameter count by 16.8% (from 0.513M to 0.427M) while achieving competitive performance comparable to the state-of-the-art on the PESQ metric (3.373). This study sets a new benchmark for the trade-off between model size and speech quality in ultra-lightweight speech enhancement.
@@ 3.1 Amplitude-Aware Linear Attention (MALA)
The MET module in the original MUSE model uses Taylor expansion to approximate Softmax attention. While this reduces complexity, the Taylor approximation (T-MSA) destroys the influence of the Query vector's norm on the attention distribution. In Softmax attention, a Query with a larger norm produces a sharper (more selective) distribution; in standard linear attention, this property is lost, causing the distribution to average out.
@@ 3.2 Inception Depthwise Convolution Embedding (IDConv)
In the time-frequency spectrogram of speech, features (formants) often exhibit specific shapes, such as vowel formants appearing as continuous horizontal stripes, while transient noise may appear as vertical wideband impulses. Traditional square kernels are inefficient at capturing these anisotropic features. MUSE uses deformable convolution to solve this, but at a high computational cost.
@@ 4.3 Comparison with State-of-the-Art Methods
As shown in the table, IMSE achieves a PESQ score of 3.373 with only 0.427M parameters.
@@
    Comparison with baseline MUSE: IMSE reduces the parameter count by 16.8% (0.427M vs. 0.513M) while maintaining a high PESQ score of 3.373 (comparable to 3.370), with slight gains in CSIG and COVL. This proves our proposed MALA and IDConv modules are not only lighter but also possess more efficient feature representation capabilities, enabling high-quality reconstruction with fewer parameters.
@@ 4.4 Ablation Study
  • Combined Effect: IMSE (Row 4) combines both. Although its PESQ (3.373) is slightly lower than the version with only IDConv, its parameter count reaches an extreme low of 0.427M. This represents an optimal trade-off under severe resource constraints, making it highly suitable for on-device applications where model size is critical.
@@ 5 Conclusion
This paper proposed IMSE, an ultra-lightweight speech enhancement network. Targeting the efficiency bottlenecks in the existing SOTA model MUSE, we utilized Inception Depthwise Convolution (IDConv) to efficiently capture multi-scale spectrogram features and Amplitude-Aware Linear Attention (MALA) to resolve the flaws of traditional linear attention. Experimental results strongly demonstrate that IMSE achieves an optimal balance between lightweight design and high performance, achieving a PESQ score of 3.373 with an extremely low parameter count of 0.427M. Future work will focus on deploying this model on actual low-power chips and exploring its application in real-time speech communication.

### v2

@@ Abstract
Achieving a balance between lightweight design and high performance remains a significant challenge for speech enhancement (SE) tasks on resource-constrained devices. Existing state-of-the-art methods, such as MUSE, have established a strong baseline with only 0.51M parameters by introducing a Multi-path Enhanced Taylor (MET) transformer and Deformable Embedding (DE). However, MUSE suffers from structural redundancy due to its approximate-then-compensate attention mechanism and computationally expensive deformable offsets. We propose IMSE to resolve these bottlenecks via: 1) Amplitude-Aware Linear Attention (MALA), which fundamentally rectifies magnitude loss in linear attention, eliminating the need for auxiliary compensation branches; and 2) Inception Depthwise Convolution (IDConv), which efficiently captures the anisotropic features of spectrograms (e.g., harmonic strips) using decomposed static kernels instead of heavy dynamic deformations. Extensive experiments on the VoiceBank+DEMAND dataset demonstrate that IMSE significantly reduces the parameter count by 16.8% (from 0.513M to 0.427M) while achieving superior performance compared to the MUSE baseline. Specifically, it achieves a PESQ score of 3.399, setting a new state-of-the-art benchmark for ultra-lightweight speech enhancement models.
@@ 3.1 Amplitude-Aware Linear Attention (MALA)
The MET module approximates Softmax via Taylor expansion but sacrifices the Query's magnitude information, resulting in over-smoothed attention. Consequently, MUSE requires a redundant branch to compensate. We instead adopt MALA, which mathematically rectifies this flaw by reintroducing magnitude via a division-based normalization. This restores the sharpness of Softmax attention at $O(N)$ complexity without structural redundancy.
@@ 3.2 Inception Depthwise Convolution Embedding (IDConv)
Speech spectrograms exhibit strong anisotropy—continuous temporal formants appear as horizontal strips, while wideband noise appears as vertical strips. Unlike MUSE's DSDCN which expends resources learning dynamic shapes, IDConv explicitly targets these physics-driven patterns using decomposed strip convolutions ($1\times K,K\times 1$). This captures acoustic textures with a static, lightweight structure, avoiding the latency of bilinear interpolation.
@@ 4.3 Comparison with State-of-the-Art Methods
As shown in Table 1, IMSE achieves a PESQ score of 3.399 with only 0.427M parameters.
@@
    Comparison with baseline MUSE: IMSE not only reduces the parameter count by 16.8% (0.427M vs. 0.513M) but also improves the speech quality across all metrics. Notably, PESQ increases from 3.370 to 3.399, and CSIG/COVL see significant gains (4.67 vs 4.63, 4.14 vs 4.10). This demonstrates that our lightweight design extracts features more effectively than the redundant structures in MUSE.
@@ 4.4 Ablation Study
  • Combined Effect: IMSE (Row 4) integrates both MALA and IDConv. Remarkably, this combination achieves the highest performance across all metrics, with a PESQ of 3.399, surpassing both the baseline and the single-module variants (e.g., 3.392 for IDConv-only).
    This result reveals a strong synergistic effect: MALA captures global amplitude-aware dependencies, while IDConv meticulously extracts anisotropic local spectral features. The combination allows the model to achieve peak performance with the lowest parameter count (0.427M), proving that efficiency and high-quality reconstruction can be achieved simultaneously without compromise.
@@ 5 Conclusion
This paper proposed IMSE, an ultra-lightweight speech enhancement network. Targeting the efficiency bottlenecks in the existing SOTA model MUSE, we utilized Inception Depthwise Convolution (IDConv) to efficiently capture multi-scale spectrogram features and Amplitude-Aware Linear Attention (MALA) to resolve the flaws of traditional linear attention. Experimental results strongly demonstrate that IMSE achieves an optimal balance between lightweight design and high performance. It achieves a PESQ score of 3.399 with an extremely low parameter count of 0.427M, establishing a new benchmark for efficiency. Future work will focus on deploying this model on actual low-power chips and exploring its application in real-time speech communication.